OpenStudy (barah101):

how do i find y^2-2y=3

jabez177 (jabez177):

You mean how do you find Y?

jabez177 (jabez177):

Because that seems like a perfect equation right there.

OpenStudy (tkhunny):

 I hope you mean, how can I find the values that can be substituted for y to make the onditional statement (equation) true! It is not a perfect squareequation. To do that, both the y^2 term and the constant term have to be perfect squares. That doesn't mean you can't "Complete the Square" and make it look like that.

OpenStudy (dayakar):

add 1 both sides to make a perfect square y^2 -2.y.1 + 1^2 = 3+1 (y - 1)^2 = 4 (y-1)^2 =2^2 can u solve for y